A második valószínűségszámítás dolgozat eredménye

1.  2.  3.  4.  5.  6.

diákok:

Bordi István  ±  +  +  +  −  +
Boros László  +  +  +  +  −  +

Böszörményi Zsolt

Csipkés László  +  +  +  +  ±  +
Ferenczik Zoltán  +  −  +  +  ±  +
Fogas János  −  −  −  −  −  −
Havasi Balázs  +  +  +  −  ±  +
Hennel Zoltán  −  −  +  ±  −  −
Illés József  +  +  +  +  ±  +
Jágó Jácint  +  +  −  +  −  +
Kiss Róbert  +  +  +  ±  −  +
Lászlók Zsolt  +  +  +  +  −  +
Lázár Gábor  +  +  +  −±  −  +
Mezei Tamás  +  +  +  +  −  +
Molnár Adrienn  +  +  +  +  −  +
Molnár Zoltán  ±  ±  +  ±  +  +
Nagy András  +  ±  ±  +  ±  +
Paper Andrea  +  ±  +  +  −  +
Pelejtei Gergely  ±  ±  +  ±  −  −
Szalontai Gábor  +  +  +  +  ±  +
Tordai Jácint  +  +  +  +  ±  +
Ujjobbágy Attila  +  +  ±  +  +  +
Vágner József  +  ±  +  −  −  −
Virág Gabriella  +  +  +  +  −  +

Zádor István